### v1

@@ 2.1 Computational Models
We begin with some unconventional terminology, particular to our paper. The class of functions $G_{\mathcal{D}}$ emerges from the interaction between abstract machines and symbolic environments that house the input-output domain ${\mathcal{D}}$. We define an abstract machine as a computable dynamical system with a discrete-time evolution. When one supplies an abstract machine with a formalized input-setting and output-reading convention, then the resulting system becomes a computational model on ${\mathcal{D}}$. Let ${\mathcal{M}}_{\mathcal{D}}$ denote the set of models computing the general recursive functions in $G_{\mathcal{D}}$. For a model $M\in{\mathcal{M}}_{\mathcal{D}}$, we use $D_{M}\subseteq{\mathcal{D}}$ to denote the set of inputs for which $M$ halts.
@@ 3 Learning from Input-Output Observations
In this original formulation of learning in the limit, the learner ${\mathcal{L}}$ observes only inputs and outputs, with no additional information about the computation process. Thus, ${\alpha}(\cdot,\cdot)\equiv{\varepsilon}$, and the model family ${\mathcal{M}}$ contributes nothing beyond function evaluations. We refer to this setting as the $\mathbf{{\mathcal{F}}({\mathcal{M}})}$–LP.
@@
For a model class ${\mathcal{M}}\subseteq{\mathcal{M}}_{\mathcal{D}}$ over an enumerable set ${\mathcal{D}}$ and for a computable monotonically increasing function $Q:{\mathbb{N}}\times{\mathbb{N}}\to{\mathbb{N}}$, let $\mathbf{TIME}_{\mathcal{D}}^{\mathcal{M}}(Q(c,n))$ denote the subset of general recursive functions $f\in{\mathcal{F}}_{\mathcal{M}}$ for which there exists a computational model $M\in{\mathcal{M}}$ – not necessarily a TM – that computes $f$ on halting inputs $x\in{\mathcal{D}}_{f}$ of $\texttt{size}_{\mathcal{D}}(x)=n$ in $O(Q(c,n))$ steps. We define the following complexity class $\mathbf{Q}({\mathcal{M}},{\mathcal{D}})=\bigcup_{c\in{\mathbb{N}}}\mathbf{TIME}_{\mathcal{D}}^{\mathcal{M}}(Q(c,n))$.
Now, under a relaxed form of the Extended Church-Turing Thesis, we prove that the concept class $\mathbf{Q}({\mathcal{M}}_{\mathcal{D}},{\mathcal{D}})$ is learnable in the limit from input-output observations and arbitrary input sources.
@@
Assuming the $q$-ECTT (2.4) holds, for an enumerable set ${\mathcal{D}}$ and a computable monotonically increasing function $Q:{\mathbb{N}}\times{\mathbb{N}}\to{\mathbb{N}}$, there exists an algorithm which learns the $Q-$time-bounded general recursive functions over ${\mathcal{D}}$ from input-output observations and restricted inputs. Formally, there exists an algorithm solving the $\mathbf{Q}({\mathcal{M}}_{\mathcal{D}},{\mathcal{D}})$–LP.
@@
Now, even if the $q$-ECTT fails, we can still salvage an important corollary if we stick to complexity classes over Turing machines: i.e., $SS\mathbf{Q}({\mathrm{T}_{\Sigma}^{\Gamma}},\SS)$. Then, the proof of Theorem 3.3 readily extends to this setting by simply replacing the encoding $\varphi$ with the identity and letting $q(c,n)=n$. Hence, popular complexity classes like $\mathbf{P},\mathbf{NP},\mathbf{\#P},\mathbf{L},$ and $\mathbf{EXP}$ become learnable in the limit.
@@
There exists an algorithm for learning any parametrized complexity class of general recursive functions in the limit from input-output observations and restricted inputs. Formally, for any finite problem alphabet $\Sigma$, tape alphabet ${\Gamma}\supseteq\Sigma\sqcup\{{\lambda}\}$, and computable monotonically increasing function $Q:{\mathbb{N}}\times{\mathbb{N}}\to{\mathbb{N}}$, there exists an algorithm solving the $SS\mathbf{Q}({\mathrm{T}_{\Sigma}^{\Gamma}},\SS)$–LP.
@@
Theorem 4.1 (Universal Learning with TBOs).
@@ A.3 Encodings and Simulations
${{\mathcal{D}}}$${{\mathcal{D}}}$$SS{\SS}$$SS{\SS}$$\scriptstyle{f}$$\scriptstyle{\varphi}$$\scriptstyle{\varphi}$$\scriptstyle{T}$
@@ Appendix B IOO–earning Proofs
We want to show that for all $f\in\mathbf{Q}({\mathcal{M}}_{\mathcal{D}},{\mathcal{D}})$, all input sources $SSI\subseteq\SS$, and all surjective orderings $w:{\mathbb{N}}{\twoheadrightarrow}I\cap D_{f}$ of the example set $E_{f}(I\cap D_{f})=\{(x,f(x)):x\in I\cap D_{f}\}$, after some finite time $t^{\star}=t^{\star}(f,I,w)$, our learning algorithm ${\mathcal{L}}$ will converge to a correct representation of $f$. In other words, we want to show that $\forall t\geq t^{\star},\ h_{t|I\cap D_{f}}=f_{|I\cap D_{f}}$, where $h_{t}$ is the hypothesis generated by ${\mathcal{L}}$ at time $t$.
Let us fix a ground truth $f\in\mathbf{Q}({\mathcal{M}}_{\mathcal{D}},{\mathcal{D}})$, an input source $SSI\subseteq\SS$, and a surjective ordering of the inputs $w:{\mathbb{N}}{\twoheadrightarrow}I\cap D_{f}$ s.t. $I\cap D_{f}=\{w_{t}\}_{t=1}^{\infty}$. We will demonstrate how ${\mathcal{L}}$ learns $f$ in the limit.
First, since $f\in\mathbf{Q}({\mathcal{M}}_{\mathcal{D}},{\mathcal{D}})$, there exists $c_{1}\in{\mathbb{N}}$ s.t. $f\in\mathbf{TIME}_{\mathcal{D}}^{\mathcal{M}}(Q(c_{1},n))$. Hence, there exists a computational model $M\in{\mathcal{M}}_{\mathcal{D}}$ which computes $f$ in $c_{2}\cdot Q(c,n)$ steps for some constant $c_{2}\in{\mathbb{N}}$. Now, let ${\mathcal{L}}$ use some reasonable encoding $SS\varphi:{\mathcal{D}}\to\SS$ for which the $q$-ECTT kicks in. Using the $q$-ECTT, there exists a TM $\widetilde{T}\in{\mathrm{T}_{\Sigma}^{\Gamma}}$ which simulates $m$ computation steps of $M$ in at most $q(c_{3},m)$ steps for some $c_{3}\in{\mathbb{N}}$. Hence, for every $x\in I\cap D_{f}$, $\widetilde{T}$ starts with $\varphi(x)$ on the tape and halts with $\varphi(f(x))$ after $q(c_{3},c_{2}\cdot Q(c_{1},|x|))$ steps. Let $c=\max(c_{1},c_{2},c_{3})$. Then, $\widetilde{T}$ computes $f$ in $q(c,c\cdot Q(c,n))$ time.

### v2

@@ 2.1 Computational Models
We begin with some unconventional terminology, particular to our paper. The class of functions $G_{\mathcal{D}}$ emerges from the interaction between abstract machines and symbolic environments that house the input-output domain ${\mathcal{D}}$. We define an abstract machine as a computable dynamical system with a deterministic discrete-time evolution. When one supplies an abstract machine with a formalized input-setting and output-reading convention, then the resulting system becomes a computational model on ${\mathcal{D}}$. Let ${\mathcal{M}}_{\mathcal{D}}$ denote the set of models computing the general recursive functions in $G_{\mathcal{D}}$. For a model $M\in{\mathcal{M}}_{\mathcal{D}}$, we use $D_{M}\subseteq{\mathcal{D}}$ to denote the set of inputs for which $M$ halts.
@@ 3 Learning from Input-Output Observations
In this original formulation of learning in the limit, the learner ${\mathcal{L}}$ observes only inputs and outputs, with no additional information about the computation process. Thus, ${\alpha}(\cdot,\cdot)\equiv{\varepsilon}$, and the model family ${\mathcal{M}}$ contributes nothing beyond function evaluations. We refer to this setting as the $\mathbf{{\mathcal{F}}_{\mathcal{M}}}$–LP.
@@
For a model class ${\mathcal{M}}\subseteq{\mathcal{M}}_{\mathcal{D}}$ over an enumerable set ${\mathcal{D}}$ and for a computable monotonically increasing function $Q:{\mathbb{N}}\times{\mathbb{N}}\to{\mathbb{N}}$, let $\mathbf{TIME}_{\mathcal{M}}(Q(c,n))$ denote the subset of general recursive functions $f\in{\mathcal{F}}_{\mathcal{M}}$ for which there exists a computational model $M\in{\mathcal{M}}$ – not necessarily a TM – that computes $f$ on halting inputs $x\in{\mathcal{D}}_{f}$ of $\texttt{size}_{\mathcal{D}}(x)=n$ in $O(Q(c,n))$ steps. We define the following complexity class $\mathbf{Q}({\mathcal{M}})=\bigcup_{c\in{\mathbb{N}}}\mathbf{TIME}_{\mathcal{M}}(Q(c,n))$.
Now, under a relaxed form of the Extended Church-Turing Thesis, we prove that the concept class $\mathbf{Q}({\mathcal{M}}_{\mathcal{D}})$ is learnable in the limit from input-output observations and arbitrary input sources.
@@
Assuming the $q$-ECTT (2.4) holds, for an enumerable set ${\mathcal{D}}$ and a computable monotonically increasing function $Q:{\mathbb{N}}\times{\mathbb{N}}\to{\mathbb{N}}$, there exists an algorithm which learns the $Q-$time-bounded general recursive functions over ${\mathcal{D}}$ from input-output observations and restricted inputs. Formally, there exists an algorithm solving the $\mathbf{Q}({\mathcal{M}}_{\mathcal{D}})$–LP.
@@
Now, even if the $q$-ECTT fails, we can still salvage an important corollary if we stick to complexity classes over Turing machines: i.e., $\mathbf{Q}({\mathrm{T}_{\Sigma}^{\Gamma}})$. Then, the proof of Theorem 3.3 readily extends to this setting by simply replacing the encoding $\varphi$ with the identity and letting $q(c,n)=n$. Hence, popular complexity classes like $\mathbf{P},\mathbf{NP},\mathbf{\#P},\mathbf{L},$ and $\mathbf{EXP}$ become learnable in the limit.
@@
There exists an algorithm for learning any parametrized complexity class of general recursive functions in the limit from input-output observations and restricted inputs. Formally, for any finite problem alphabet $\Sigma$, tape alphabet ${\Gamma}\supseteq\Sigma\sqcup\{{\lambda}\}$, and computable monotonically increasing function $Q:{\mathbb{N}}\times{\mathbb{N}}\to{\mathbb{N}}$, there exists an algorithm solving the $\mathbf{Q}({\mathrm{T}_{\Sigma}^{\Gamma}})$–LP.
@@
Theorem 4.1 (Universal TBO-Learning).
@@ A.3 Encodings and Simulations
${{\mathcal{D}}}$${{\mathcal{D}}}$$SS{\SS}$$SS{\SS}$$\scriptstyle{M}$$\scriptstyle{\varphi}$$\scriptstyle{\varphi}$$\scriptstyle{T}$
@@ Appendix B IOO–earning Proofs
We want to show that for all $f\in\mathbf{Q}({\mathcal{M}}_{\mathcal{D}})$, all input sources $SSI\subseteq\SS$, and all surjective orderings $w:{\mathbb{N}}{\twoheadrightarrow}I\cap D_{f}$ of the example set $E_{f}(I\cap D_{f})=\{(x,f(x)):x\in I\cap D_{f}\}$, after some finite time $t^{\star}=t^{\star}(f,I,w)$, our learning algorithm ${\mathcal{L}}$ will converge to a correct representation of $f$. In other words, we want to show that $\forall t\geq t^{\star},\ h_{t|I\cap D_{f}}=f_{|I\cap D_{f}}$, where $h_{t}$ is the hypothesis generated by ${\mathcal{L}}$ at time $t$.
Let us fix a ground truth $f\in\mathbf{Q}({\mathcal{M}}_{\mathcal{D}})$, an input source $SSI\subseteq\SS$, and a surjective ordering of the inputs $w:{\mathbb{N}}{\twoheadrightarrow}I\cap D_{f}$ s.t. $I\cap D_{f}=\{w_{t}\}_{t=1}^{\infty}$. We will demonstrate how ${\mathcal{L}}$ learns $f$ in the limit.
First, since $f\in\mathbf{Q}({\mathcal{M}}_{\mathcal{D}})$, there exists $c_{1}\in{\mathbb{N}}$ s.t. $f\in\mathbf{TIME}_{\mathcal{M}}(Q(c_{1},n))$. Hence, there exists a computational model $M\in{\mathcal{M}}_{\mathcal{D}}$ which computes $f$ in $c_{2}\cdot Q(c,n)$ steps for some constant $c_{2}\in{\mathbb{N}}$. Now, let ${\mathcal{L}}$ use some reasonable encoding $SS\varphi:{\mathcal{D}}\to\SS$ for which the $q$-ECTT kicks in. Using the $q$-ECTT, there exists a TM $\widetilde{T}\in{\mathrm{T}_{\Sigma}^{\Gamma}}$ which simulates $m$ computation steps of $M$ in at most $q(c_{3},m)$ steps for some $c_{3}\in{\mathbb{N}}$. Hence, for every $x\in I\cap D_{f}$, $\widetilde{T}$ starts with $\varphi(x)$ on the tape and halts with $\varphi(f(x))$ after $q(c_{3},c_{2}\cdot Q(c_{1},|x|))$ steps. Let $c=\max(c_{1},c_{2},c_{3})$. Then, $\widetilde{T}$ computes $f$ in $q(c,c\cdot Q(c,n))$ time.